OpenStudy (anonymous):

Can someone help me solve x2 + 9x +3 = 0?

OpenStudy (anonymous):

well you can factor it

OpenStudy (zzr0ck3r):

(x+4.5)^2=-3+ 4.5^2

OpenStudy (zzr0ck3r):

can you solve that?

OpenStudy (anonymous):

Yes, thank you. :)

OpenStudy (anonymous):

EXAMPLE FACTORING X ^2 + X - 12 = 0 (X + 4) ( X - 3 ) = 0 X = -4 , 3

OpenStudy (zzr0ck3r):

@iforgot this does not factor.

OpenStudy (anonymous):

that was an example

OpenStudy (anonymous):

you can try quadratic formula

OpenStudy (zzr0ck3r):

yes there are many ways to solve, I was just letting you know it does not factor.

OpenStudy (anonymous):

the sure way most tedious also to get the answer

OpenStudy (anonymous):

o okies :P

OpenStudy (zzr0ck3r):

completing the square is the only way to go:)

OpenStudy (anonymous):

yeah i usually give esamples rarely do i go work with ppl on problems